OpenStudy (anonymous):

Will someone quickly explain to me how to simplify this rational expression? Thanks!

OpenStudy (anonymous):

\[\frac{2x^2 + 6x }{ 10x^3 }\]

OpenStudy (anonymous):

id start by factoring. what can you pull out of both the numerator and the denominator.

OpenStudy (anonymous):

The number 2?

OpenStudy (anonymous):

2x, sorry.

OpenStudy (anonymous):

yep, you can also cancel something as well.

OpenStudy (anonymous):

the most important thing about canceling is that the term u are canceling must be in each term in the numerator and denominator. for example.. x+3/x^2 you cannot cancel an x because there is no x with the 3

OpenStudy (anonymous):

so same here.. you cannot cancel x^2 because only a single power x term is in the numerator's 2nd term.

OpenStudy (anonymous):

we.. if you pull out 2x, you would be left with \[\frac{ x+3 }{ 5x^2 }\]

OpenStudy (anonymous):

NO WAIT! \[\frac{ x + 3 }{ 5x^2 }\] Ach! You typed it in before me. Oh well. Thank you, you have been very helpful!

OpenStudy (anonymous):

great! ill be around for a few hours. let me know if you need more help.